OpenStudy (anonymous):

What is the exact value of the expression 8 sqrt 15 + 5 sqrt 5 +sqrt 15

OpenStudy (ghazi):

\[8 \sqrt 15+ 5 \sqrt 5 + \sqrt 15= 9 \sqrt{15}+5 \sqrt 5\]

OpenStudy (ghazi):

can you do it further?

OpenStudy (ghazi):

15=5*3

OpenStudy (anonymous):

oh so the answer is sqrt 35

OpenStudy (ghazi):

no

OpenStudy (anonymous):

no u cant do it further

OpenStudy (anonymous):

thats the answer

OpenStudy (anonymous):

the thing u wrote is right

OpenStudy (ghazi):

i can do it further

OpenStudy (ghazi):

\[9 \sqrt 15+ 5 \sqrt 5= 9\sqrt3 \sqrt5+ 5 \sqrt5= (9 \sqrt3+ 5) \sqrt5\]

OpenStudy (anonymous):

so the answer is 8 sqrt 15 + 5 sqrt 5

OpenStudy (ghazi):

no the answer would be either 9 sqrt 15+ 5 sqrt 5 or the one i have written above

OpenStudy (ghazi):

@zahabaha clear?

OpenStudy (anonymous):

thats wat i said

OpenStudy (ghazi):

you said 8 sqrt 15 + 5 sqrt 5

OpenStudy (anonymous):

i said wat u wrote is right it cant go further

OpenStudy (anonymous):

read above

OpenStudy (ghazi):

i did it further...

OpenStudy (ghazi):

well i guess it is solved isn't it?